I need to know how to solve the following problem:
2X+4(18-X)=58

Answer by stanbon(75887)

I need to know how to solve the following problem:
2X+4(18-X)=58
---
2x + 72 - 4x = 58
-2x = -14
x = 7
======
Cheers,
Stan H.